# Aspects of the commuting graph [^1] To the memory of Richard Parker

### v3

@@ 7.1 Isoclinism
    Check that there is one equivalence class of size $p$, consisting of the dominating vertices, and the other equivalence classes all have size $p^{2}-p$.
  • (d)
    Now we can compute the graph $\Gamma$ by choosing $|C|/p$ vertices from each equivalence class $C$. This gives a graph with $p^{2a}$ vertices.
It remains to recognise the symplectic graph. There are two possible strategies we could adapt. The first is to build, from the symplectic graph, a vector space with a bilinear form so that we have explicitly identified the unknown graph. We were able to produce a polynomial-time algorithm for this, which we hope to publish elsewhere; but in the interests of brevity we adopt a different strategy; our algorithm will simply decide if our unknown graph is isomorphic to a symplectic graph. The strategy is to go from the vector space to the projective space and use the Buekenhout–Shult theorem [16] to recognise the graph. For reference, we state the theorem in the finite case.
@@
Let $(P,L)$ be a finite point-line geometry, with $P$ and $L$ both non-empty. Assume
  1. (a)
    two points lie on at most one line;
  2. (b)
    no point is collinear with all others;
  3. (c)
    if a point $p$ is not on a line $l$, then $p$ is joined to one or all points of $l$.
Then either $(P,L)$ is a generalized quadrangle, or it is the geometry of points and totally isotropic (or totally singular) lines in a classical polar space over a finite field.
We refer to the paper for the undefined terms in this theorem.
Now we proceed in a number of steps.
Step 1
The case $a=1$. A non-abelian group of order $p^{3}$ is the union of $p+1$ maximal abelian subgroups of order $p^{2}$ intersecting pairwise in the identity; so our original graph is the commuting graph of a non-abelian group of order $p^{3}$ if and only if the graph $\Gamma$ consists of $p+1$ pairwise disjoint cliques of size $p-1$ together with a dominating vertex.
Step 2
Moving to projective space. For this we remove the dominating vertex (which would correspond to the zero vector) and choose one from each equivalence class of size $p-1$ for the remaining vertices. We obtain a graph on $(p^{2a}-1)/(p-1)$ vertices.
Step 3
Construct the lines. If $u$ and $v$ are adjacent vertices of $\Gamma$, the line joining them is
$L(u,v)=\{u,v\}\cup\{w:\bar{N}(u)\cap\bar{N}(v)\subseteq\bar{N}(w)\}.$
If a line contains fewer than $p+1$ points, the algorithm terminates and returns false.
Step 4
Check the Buekenhout–Shult hypotheses. Return false if they fail.
We are left with the two possible conclusions of Buekenhout and Shult. The first of these is that the geometry is a generalized quadrangle. (This will only occur if $a=2$.) In this case, the points of the quadrangle are all the points of the $3$-dimensional projective space over the $p$-element field $F$. The closed neighbourhoods are the planes of the projective space. It is clear that the closed neighbourhood of a vertex contains the line through any two of its points. So it suffices to show that all closed neighbourhoods have size $p^{2}+p+1$ to identify the projective space. If so, then the map $v\leftrightarrow\bar{N}(v)$ is a null polarity of the projective space, and so defines a symplectic polarity; the algorithm can return true.
In the other case, there is still one problem to resolve. Inspecting the list of classical polar spaces, we see that there is another with the same parameters as the symplectic space on a $2k$-dimensional vector space: the orthogonal space, defined by a non-singular quadratic form on a $(2k+1)$-dimensional vector space. In characteristic $2$, they are isomorphic, so there is no more to do. In odd characteristic, the orthogonal space contains a $3\times 3$ grid. Indeed, over the $p$-element field $F$, it contains a $(p+1)\times(p+1)$ grid, the quadric defined by a hyperbolic form (one of Witt index $2$) on a $4$-dimensional vector space. (This is also realised as a regulus in projective $3$-space together with the opposite regulus.) But the symplectic space contains no such grid. For we can take a $4$-cycle with vertices spanned by the vectors $e_{1},e_{2},f_{1},f_{2}$ (where $b(e_{i},e_{j})=b(f_{i},f_{j})=0$ and $b(e_{i},f_{j})=\delta_{ij}$). Choose arbitrary points on the edges of the cycle so that opposite points are collinear: say the points spanned by $e_{1}+ae_{2}$, $f_{2}-af_{1}$, $e_{2}+bf_{1}$, $e_{1}+bf_{2}$. The determinant of their coefficients is $2ab\neq 0$, so the two lines are skew.
Thus we can simply check whether such a grid exists. Transitivity of the group of the symplectic space on $4$-cycles shows that only one $4$-cycle needs to be checked.
We leave to the reader the verification that all these tests can be done in polynomial time.
@@ 8 A quasipolynomial time algorithm in the general case
Assuming CFSG (the classification of finite simple groups), in this section we present a $2^{O(\log^{3}n)}$ time algorithm for checking if a given $n$-vertex graph is the commuting graph of some $n$-element group and, if so, labeling the vertices of the graph by the group elements consistent with all the commuting pairs.
@@ 8.1 Short descriptions for finite simple groups assuming CFSG
As there are only finite many sporadic finite simple groups, for the algorithm $\mathcal{D}$ they can be trivially dealt with. The cyclic groups of prime order are also trivial.
@@
Proof 7
As $G=K/N$ is simple it can be generated by two elements $Ng_{1},Ng_{2}\in K/N$ where $g_{1},g_{2}\in K$. Let $\mathop{\mathrm{desc}}(G)$ be the indices of these two elements in $H$ along with $\mathop{\mathrm{desc}}(H)$. (By “indices” here we mean the positions of these elements in the description.) Clearly, it is of size $O(\log|H|)$. The claimed algorithm will first use $\mathop{\mathrm{desc}}(H)$ to compute the multiplication table for $H$. It can enumerate $H$ and use the indices to identify $g_{1}$ and $g_{2}$. Using the multiplication table for $H$ it can generate the subgroup $K=\langle g_{1},g_{2}\rangle$, listing down all its elements in time polynomial in $|H|$. Next, the algorithm can identify the normal subgroup $N$ (as given in the hypothesis) in time polynomial in $|H|$ and obtain the multiplication table for $G=K/N$$\qed$

### v4

@@ 7.1 Recognizing the symplectic graph
    Then notice that the subspace $\{\bar{x}_{1},\bar{y}_{1}\}^{\perp}$ is a $22r-$ 2 dimensional symplectic space $V^{\prime}$ (w.r.t. the same symplectic form). We can continue with the basis construction, recursively applying the above two steps to $V^{\prime}$.
Now, suppose $X=(V,E)$ is the graph given as input on $p^{2r}$ vertices. If $X$ is indeed the symplectic graph of $(V,2r,p)$ for some symplectic form $\langle$,$\rangle$ , ⟩, we will describe a polynomial-time recursive procedure that recovers a bijection from $V$ to $V(2r,p)$ along with a symplectic form $\langle$,$\rangle$ , ⟩ as witness for $X$ being the associated symplectic graph. We will then argue that if $X$ is not a symplectic graph then this procedure detects that.
Essentially, the problem is to efficiently simulate the above two-step basis construction procedure given only the graph $X=(V,E)$ as input. As the symplectic space is non-degenerate, the zero vector 00 is the only vertex in $X$ adjacent to all others and is easily identified. Let $x_{1}$ be any other vertex in $V$ and we choose a non-zero vector $\bar{x}_{1}\in V(2r,p)$ as its image. Next, we will choose $y_{1}$ as any vertex not adjacent to $x_{1}$ in the graph $X$ and choose a non-zero vector $\bar{y}_{1}\in V(2r,p)$ as its image. Let
$V^{\prime}=\{v\in V\mid(v,x_{1})\in E\text{ and }(v,y_{1})\in E\}.$
That is, $V^{\prime}$ is the common neighborhood of $x_{1}$ and $y_{1}$. We can easily identify $V^{\prime}$ in $X$. Let $X^{\prime}$ be the subgraph of $X$ induced by the vertex subset $V^{\prime}$. We have the following simple observation.
Claim 7.7
If $X$ is the symplectic graph of a $2r$-dimensional symplectic space over $\mathbb{F}_{p}$ then the graph $X^{\prime}$ is the symplectic graph of a symplectic space of dimension $22r-$ 2 over $\mathbb{F}_{p}$.
The algorithm will now recursively check that $X^{\prime}$ is indeed the symplectic graph of a $22r-$ 2-dimensional symplectic space $V(2r-2,p)$ over $\mathbb{F}_{p}$. If so, it will compute a labeling of the vertices of $V^{\prime}$ by linear combinations $\sum_{i=2}^{r}(\alpha_{i}\bar{x}_{i}+\beta_{i}\bar{y}_{i}),\alpha_{i},\beta_{i}\in\mathbb{F}_{p}$ that is consistent with the orthogonality relation of a symplectic form $\langle$,$\rangle$ , ⟩. Here the basis vectors computed by the algorithm are $\bar{x}_{i},\bar{y}_{i},2\leq i\leq r$. At the end of the recursive call, the algorithm computes a bijection from $V^{\prime}$ to $V(2r-2,p)$ that labels each vertex in $V^{\prime}$ by a distinct linear combination of the basis vectors.
Now, the remaining task for the algorithm is to find a labeling of the vertices in $V\setminus V^{\prime}$ consistent with a symplectic form on $V(2r,p)$, assuming $X$ is a symplectic graph.
For each vertex $v$ we will denote its vector label by $\bar{v}$. For a labeled subset $S$ of vertices of $V$ we will use $\bar{S}\subseteq V(2r,p)$ to denote the set of vectors labeling vertices in $S$.
Claim 7.8
A vertex $v\in V\setminus V^{\prime}$ can be labeled by a non-zero vector $\alpha\bar{x}_{1}+\beta\bar{y}_{1}$ in $V(2r,p)$, $\alpha,\beta\in\mathbb{F}_{p}$, if and only if $(v,u)\in E$ for all $u\in V^{\prime}$.
Proof of Claim. Consider the symplectic graph of $V(2r,p)$. Let $V(2r-2,p)$ denote the subspace spanned by $\bar{x}_{i},\bar{y}_{i},2\leq i\leq n$. Clearly, every vector of the form $\alpha\bar{x}_{1}+\beta\bar{y}_{1},\alpha,\beta\in\mathbb{F}_{p}$ is orthogonal to each vector in $V(2r-2,p)$. Conversely, consider a vector $\alpha\bar{x}_{1}+\beta\bar{y}_{1}+v\in V(2r,p)$, where $v\in V(2r-2,p)$ is non-zero. Since $V(2r-2,p)$ is non-degenerate, there is a $u\in V(2r-2,p)$ such that $\langle v,u\rangle\neq 0$ which implies $\langle\alpha\bar{x}_{1}+\beta\bar{y}_{1}+v,u\rangle=\langle v,u\rangle\neq 0$.
Thus, the algorithm can find a subset $V_{12}$ of precisely $p^{2}-1$ many such vertices in $V$ that are adjacent to all of $V^{\prime}$, of which we have already labeled two vertices as $\bar{x}_{1}$ and $\bar{y}_{1}$. We analyze the subgraph $X_{12}$ of $X$ induced by $V_{12}$.
Claim 7.9
The subgraph $X_{12}$ induced by $V_{12}$ consists of a disjoint union of $p+1$ many $(p-1)$-size cliques, $C_{1},C_{2},\ldots,C_{p+1}$. Furthermore, the labeling procedure must label the vertices of each $C_{i}$ from the set $\{\alpha\bar{u}\mid\alpha\in\mathbb{F}^{*}_{p}\}$ of $p-1$ vectors, for some non-zero vector $\bar{u}\in\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$.
Proof of Claim. It is clear from the definition of $V_{12}$ that the vertices in it must be labeled by nonzero vectors in $\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$. The proof follows easily by observing that a non-zero vector $\alpha\bar{x}_{1}+\beta\bar{y}_{1}$ is orthogonal only to itself and its scalar multiples and to no other vectors in $\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$.
By the above claim, we can consistently label each vertex $u\in V_{12}$ by a vector $\bar{u}\in\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$ assuming $X$ is a symplectic graph.
Claim 7.10
If $X=(V,E)$ is the symplectic graph of the symplectic space $V(2r,p)$ and $V^{\prime}$ corresponds to the symplectic subspace $V(2r-2,p)$ spanned by $\{\bar{x}_{i},\bar{y}_{i}\mid 2\leq i\leq r\}$ then for any vertex $u\in V_{12}$ the vertex subset
$V[u]=\{v\in V\setminus V^{\prime}\mid(v,u)\in E\}$
has to be labeled by the vectors
$\{\alpha\bar{u}+\bar{v}\mid\alpha\neq 0,\bar{v}\in\mathop{\mathrm{Span}}(\cup_{i=2}^{r}\{\bar{x}_{i},\bar{y}_{i}\})\}$
of $V(2r,p)$.
Proof of Claim. Suppose $v\in V\setminus V^{\prime}$ is labeled by the vector $\bar{a}+\bar{b}$, where $\bar{a}\in\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$ and $\bar{b}\in V(2r-2,p)$. As $v\in V[u]$ we have
$\langle\bar{a}+\bar{b},\bar{u}\rangle=0.$
As $\langle\bar{b},\bar{u}\rangle=0$ it implies that $\langle\bar{a},\bar{u}\rangle=0$ which means $\bar{a}=\alpha\bar{u}$ for some non-zero scalar $\alpha$. Conversely, any vertex labeled by $\alpha\bar{u}+\bar{b}$ for $\bar{b}\in V(2r-2,p)$ has to lie in $V[u]$ if the labeling is consistent with the symplectic graph $X$.
It follows that $|V[u]|=(p-1)\cdot p^{2r-2}$. Furthermore, if vertices $u$ and $v$ are in the same $(p-1)$-clique in the graph $X_{12}$ then $V[u]=V[v]$. As $(p+1)(p-1)\cdot p^{2r-2}=p^{2r}-p^{2r-2}=|V\setminus V^{\prime}|$, it follows that the different $V[u]$ are all pairwise disjoint and form a partition of $V\setminus V^{\prime}$. Consequently, we can identify the $p+1$ many equisized vertex subsets $V[u]$ which form a partition of $V\setminus V^{\prime}$.
Labelling vertices in $V[u]$
By Claim 7.10, the vertices of $V[u]$ have to be labeled by vectors from $\{\alpha\bar{u}+\bar{v}\mid v\in V^{\prime},\alpha\neq 0\}$. We will examine the adjacencies in the subgraph induced by $V[u]$ and also the adjacencies across the different $V[u]$ and $V[u^{\prime}]$ to determine a consistent labeling. Let $X[u]$ denote the subgraph of $X$ induced by $V[u]$.
For distinct vertices $v_{1},v_{2}\in V^{\prime}$ notice that
$\langle\bar{v}_{1},\bar{v}_{2}\rangle=\langle\alpha\bar{u}+\bar{v}_{1},\beta\bar{u}+\bar{v}_{2}\rangle.$ (7)
Thus, in any consistent labeling of $V[u]$ suppose $S_{u,v_{i}}\subset V[u]$ are the $p-1$ vertices labeled by $\{\alpha\bar{u}+\bar{v}_{i}\mid\alpha\neq 0\}$ for $i\in\{1,2\}$. Then Equation 7 implies that in the subgraph $X[u]$ the bipartite graph induced by the subsets $S_{u,v_{1}}$ and $S_{u,v_{2}}$ is either complete or empty.
Furthermore, $\langle\alpha\bar{u}+\bar{v},\beta\bar{u}+\bar{v}\rangle=0$ for all non-zero $\alpha,\beta$. Hence each $S_{u,v},v\in V^{\prime}$ induces a $(p-1)$-clique.
How do we identify each $S_{u,v}$? Suppose $v_{1},v_{2}\in V^{\prime}$ are distinct vertices. In $V^{\prime}$ (which is already labeled by $V(2r-2,p)$) we can find a vertex $v_{3}$ such that $\langle v_{1},v_{3}\rangle\neq 0$ and $\langle v_{2},v_{3}\rangle=0$. Hence each $S_{u,v}$ can be identified because
  • (a)
    All $p-1$ vertices in $S_{u,v}$ are adjacent to each other. They all have identical neighborhoods in $X[u]$ by the above argument.
  • (b)
    Every vertex in $V[u]\setminus S_{u,v}$ has a different neighborhood in $X[u]$ than the vertices in $S_{u,v}$.
Putting it together, in summary, we have shown the following.
Claim 7.11
In polynomial time we can uniquely identify all the subsets $S_{u,v},v\in V^{\prime}$, each of size $p-1$, forming a partition of $V[u]$, for each $u\in V_{12}$.
Thus, we have identified the subsets $S_{u,v}\subset V[u]$ for $v\in V^{\prime}$. We now need to label the elements of $S_{u,v}$ by vectors $\{\alpha\bar{u}+\bar{v}\mid\alpha\in\mathbb{F}^{*}_{p}\}$.
Claim 7.12
Assuming $X$ is a symplectic graph, for each $v\in V^{\prime}$ and $u\in V_{12}$, we can label the $p-1$ vertices in $S_{u,v}$ by the $p-1$ vectors $\alpha\bar{u}+\bar{v},\alpha\in\mathbb{F}^{*}_{p}$ arbitrarily to get a labeling for $V\setminus V^{\prime}$ that is consistent with the labeling of $V^{\prime}$.
Proof of Claim. We first observe that adjacencies within $V[u]$ and between $V[u]$ and $V^{\prime}$ are consistent with such a labeling by the arguments preceding the claim.
We now examine the adjacencies across two disjoint $V[u]$ and $V[u^{\prime}]$. Consider $S_{u,v}\subset V[u]$ and $S_{u^{\prime},v^{\prime}}\subset V[u^{\prime}]$. There is an edge between these subsets if and only if for some non-zero scalars $\alpha$ and $\beta$ we have $\langle\alpha\bar{u}+\bar{v},\beta\bar{u^{\prime}}+\bar{v^{\prime}}\rangle=0$. Now, $\langle\bar{u},\bar{u^{\prime}}\rangle\neq 0$. Hence we must have
$\alpha\beta=\frac{-\langle\bar{v},\bar{v^{\prime}}\rangle}{\langle\bar{u},\bar{u^{\prime}}\rangle}.$
Since $\alpha$ and $\beta$ are non-zero scalars, it follows that $\langle\bar{v},\bar{v^{\prime}}\rangle\neq 0$. It follows that $v$ and $v^{\prime}$ being non-adjacent in the graph induced by $V^{\prime}$ is a necessary condition for $S_{v}$ and $S_{v^{\prime}}$ to have any edges between them. Furthermore, the above equation implies that for the vertex, say $v_{\alpha}\in S_{u,v}$ labeled $\alpha\bar{u}+\bar{v}$ there is a unique $\beta\neq 0$ such that $v_{\alpha}$ is adjacent to the vertex, say $v^{\prime}_{\beta}$ in $S_{u^{\prime},v^{\prime}}$ labeled $\beta\bar{u^{\prime}}+\bar{v^{\prime}}$. Hence, either there are no edges between $S_{v}$ and $S_{v^{\prime}}$ (if $v$ and $v^{\prime}$ are adjacent) or there is a perfect matching between $S_{v}$ and $S_{v^{\prime}}$( if $v$ and $v^{\prime}$ are non-adjacent).
Since these are all the possible edges in $V\setminus V^{\prime}$ and it is clearly consistent with any labeling of $S_{u,v}$ by the $p-1$ vectors $\alpha\bar{u}+\bar{v},\alpha\neq 0$ for all $u\in V_{12}$ and $v\in V^{\prime}$, this completes the proof of the claim.
Putting it together, clearly if $X$ is the symplectic graph of some non-degenerate symplectic form on $V(2r,p)$ the above algorithm verifies that by constructing a symplectic form consistent with the symplectic graph. The algorithm is described as a recursive procedure below. Essentially, given a graph $X=(V,E)$ with $|V|=p^{2r}$ vertices as input it will try to compute a labeling of $V$ by the $p^{2r}$ vectors in $V(2r,p)=\{\sum_{i=1}^{r}(\alpha_{i}\bar{x}_{i}+\beta_{i}\bar{y}_{i})\mid\alpha_{i},\beta_{i}\in\mathbb{F}_{p}\}$, such that the edges of $X$ are consistent with some symplectic form on $V(2r,p)$. If $X$ is a symplectic graph it will succeed. Otherwise, the procedure will fail and detect that $X$ is not symplectic.
Procedure Label$(X,2r,p)$
  1. The procedure takes as input a simple graph $X=(V,E)$ with $p^{2r}$ vertices for an odd prime $p$ and computes a labeling consistent with some symplectic form on $V(2r,p)$ if $X$ is a symplectic graph. Otherwise, it returns “fail”.
  2. (a)
    Identify the unique dominant vertex $v_{0}\in V$ and label it by 00 (the zero vector).
  3. (b)
    Pick any pair of nonadjacent vertices $x_{1},y_{1}\in V\setminus\{v_{0}\}$ and label them by the basis vectors $\bar{x}_{1}$ and $\bar{y}_{1}$ respectively.
  4. (c)
    Let $V^{\prime}$ be the set of all vertices adjacent to both $x_{1}$ and $y_{1}$. Check $|V^{\prime}|=p^{2r-2}$. Let $X^{\prime}$ be the subgraph induced by $V^{\prime}$. Recursively call Label$(X^{\prime},2r-2,p)$ and obtain a labeling of $V^{\prime}$ by vectors in $V(2r-2,p)$ defined by the basis $\bar{x}_{i},\bar{y}_{i},2\leq i\leq p$ and consistent with a symplectic form on $V(2r-2,p)$ (Claim 7.7).
  5. (d)
    Determine the $p^{2}-1$ vertices $V_{12}$ (Claim 7.8) consisting of vertices in $V\setminus V^{\prime}$ that are adjacent to all vertices in $V^{\prime}$. Determine the partition of $V_{12}$ into $p+1$ many $(p-1)$-cliques in the induced subgraph $G_{12}$ (Claim 7.9).
  6. (e)
    From the adjacencies between $V\setminus V^{\prime}$ and $V^{\prime}$ determine the partition of $V\setminus V^{\prime}$ into $p+1$ subsets $V[u]$ (Claim 7.10).
  7. (f)
    From the subgraphs $G[u]$ induced by the respective $V[u]$, determine the partition of $V[u]$ into $p-1$-subsets $S_{u,v},v\in V^{\prime}$ (Claim 7.11) .
  8. (g)
    Label the vertices in $S_{u,v},v\in V^{\prime}$ and $u\in V_{12}$ by the vectors $\alpha\bar{u}+\bar{v},\alpha\neq 0$ (Claim 7.12).
  9. (h)
    If any step above does not satisfy the stated property then output fail. Else output pass along with the complete labeling of $V$ by elements of $V(2r,p)$ as the witness that $X$ is a symplectic graph.
Running time analysis  Let $T(n)$ denote the running time for constructing a symplectic form $\langle$,$\rangle$ , ⟩ for $V(2r,p)$ consistent with $X$ as the symplectic graph, where $n=p^{2r}$ is the number of vertices in $X$. The inductive construction and the rest of the computation implies the recurrence $T(n)=O(|V|^{2})+T(n-1)$, which gives an overall cubic bound $T(n)=O(|V|^{3})=O(p^{6r+3})$ on the running time.
Putting it together, we have a polynomial-time algorithm that checks if $X$ is the symplectic graph for the symplectic space by finding out a labeling of vertices by the vectors along with a consistent symplectic form. $\qed$
In summary, we have the following.
@@
Given a graph $X=(V,E)$ with $p^{2r+1}$ vertices we can determine in polynomial time if it is the commuting graph of an extraspecial group of order $p^{2r+1}$ and, if so, label vertices by unique group elements satisfying the commuting relation.
Proof 13
By Lemma 7.3 we can obtain the graph $X_{s}=(V_{s},E_{s})$ from $X$ in polynomial time. By Theorem 7.6 we can check if $X_{s}$ is the symplectic graph of a symplectic form $\langle$,$\rangle$ , ⟩ on $V(2n,p)$ and also find the symplectic form.
From the proof of Lemma 7.3, we have the following observations.
  • (a)
    The vertex of degree $|V_{s}|-1$ in $X_{s}$ is the 00 element of $V(2n,p)$, and corresponds to $Z$ in $G$.
  • (b)
    For each vertex $x\in V_{s}$ there is a closed-twins equivalence class of size $p-1$ containing $x$. There is a corresponding subset of vertices $H_{x}\setminus Z$ in $X$ of size $p^{2}-p$, where $H_{x}$, in turn, corresponds to a subgroup of $G$ of order $p^{2}$ that contains $Z$. We have a labeling of the vertex $x$ by a linear combination $\sum_{i=1}^{n}(\alpha_{i}\bar{x}_{i}+\beta_{i}\bar{y}_{i})$, and the remaining $p-2$ vertices in the equivalence class are labeled by non-zero scalar multiples of this linear combination.
Now, using the description of the extraspecial group $G_{1}$ we can label the vertices of the clique $H_{x}\setminus Z$ with the group elements
$z^{j}\cdot\left(\prod_{i=1}^{n}x_{i}^{\alpha_{i}}.y_{i}^{\beta_{i}}\right)^{k},0\leq j\leq p-1,1\leq k\leq p-1.$
This will give us the labeling of $X$ with the elements of the extraspecial group $G_{1}$ of order $p^{2r+1}$, consistent with the commuting graph $X$.
Observe that if $X_{s}$ is not a symplectic graph then Procedure Label described in Theorem 7.6 will detect that and return fail. This completes the proof. $\qed$
@@ 8 A quasipolynomial time algorithm in the general case
Assuming CFSG (the classification of finite simple groups [41]), in this section we present a $2^{O(\log^{3}n)}$ time algorithm for checking if a given $n$-vertex graph is the commuting graph of some $n$-element group and, if so, labeling the vertices of the graph by the group elements consistent with all the commuting pairs. For CFSG in this section we will refer to the (concrete) descriptions of the finite simple groups in Wilson’s text [79].
@@ 8.1 Short descriptions for finite simple groups assuming CFSG
As there are only finite many sporadic finite simple groups, for the algorithm $\mathcal{D}$ they can be trivially dealt with. The cyclic groups of prime order are also easy to deal with for the algorithm $\mathcal{D}$ as there is a unique such subgroup of order $p$ for each prime $p$.
@@
Proof 14
The algorithm $\mathcal{D}$ will take the binary encoding $\mathop{\mathrm{bin}}(p)$ of $p$, verify that $p$ is prime using the polynomial-time primality test, and then write out the group table for the additive group of integers modulo $p$. $\qed$
@@
Proof 15
As $G=K/N$ is simple it can be generated by two elements $Ng_{1},Ng_{2}\in K/N$ where $g_{1},g_{2}\in K$. Let $\mathop{\mathrm{desc}}(G)$ be the indices of these two elements in $H$ along with $\mathop{\mathrm{desc}}(H)$. By indices here we mean the positions of these elements in the multiplication table for $H$. Clearly, $\mathop{\mathrm{desc}}(G)$ is of size $O(\log|H|)$. The claimed algorithm $\mathcal{D}$ will first use $\mathop{\mathrm{desc}}(H)$ to compute the multiplication table for $H$. It can enumerate $H$ and use the indices to identify $g_{1}$ and $g_{2}$. Using the multiplication table for $H$ it can generate the subgroup $K=\langle g_{1},g_{2}\rangle$, listing down all its elements in time polynomial in $|H|$. Next, the algorithm can identify the normal subgroup $N$ (as given in the hypothesis) in time polynomial in $|H|$ and obtain the multiplication table for $G=K/N$$\qed$